# LLM-driven Multimodal Target Volume Contouring in Radiation Oncology

### v3

@@ Results
Tuning Method Multiple text prompts - Multiple text prompts One text prompt LoRA [28] No tuning
@@
External Test #2 (N=205) Dice $\uparrow$ 0.844 (0.826-0.857) 0.737 (0.717-0.755) 0.528 (0.490-0.567) 0.832 (0.815-0.846) 0.796 (0.779-0.810) 0.829 (0.812-0.842)
@@
numeric category of each clinical data, e.g., “0301”, corresponds to “N0” for the first digit, ‘T3” for the second, “mastectomy” for third, and “right” for the last.
We further ablated our employment of introducing textual information by replacing it with various methodologies. These include utilizing a single or multiple text prompts through prompt tuning, low-rank adaptation (LoRA) fine-tuning [28] and directly employing a pre-trained LLM without tuning. As indicated in Table 3(b), our proposed text prompt tuning method consistently outperformed those using LoRA fine-tuning and no tuning strategy. Moreover, employing multiple learnable text prompts showed an improved performance compared to using a single text prompt. These results indicate that the introduced learnable text prompts were optimized to efficiently fine-tune the LLM for the target volume contouring task.
Ablation Study of Input Textual Information Components.
Figure 4: Qualitative comparison between different multimodal methods with omitted input text information components.
We further conducted ablation study by omitting each piece of input textual information and compared the difference between a competing method (Numeric Category) and our method (LLMSeg) in Fig. 4(b)-(c). Firstly, without omission as shown in Fig. 4(b), our method accurately segmented only the right breast area as the target volume for a person with T1aN0M0 cancer who underwent breast-conserving surgery. However, when the information for T stage was removed, the model confused and included some regional nodes in the target range as if considering it an upstage. This trend was similarly observed in the omission of N stage information, where the model included regional nodes as in cases with nodal metastasis like N1 or N2. Likewise, without information about laterality, the model contours on the opposite breast. On the contrary, the competing model showed inaccurate results such as contouring on the opposite breast even without omission. Moreover, regardless the presence or absence of omission, there was little change in target contouring (e.g., laterality), or target contouring changed in patterns unrelated to the omitted information (e.g., contouring on the opposite side when omitting T stage or N stage information). These results indicate that the competing model receiving clinical context in a simpler manner fail to effectively incorporate such information and perform CTV delineation unrelated to the provided information. Similarly, in another case of T1cN1M0 breast cancer in the left breast where total mastectomy was performed as shown in Fig. 4(c), when surgery information is not provided, our method confused the surgery type and showed ablation results sparing the skin and chest wall, like the case of breast-conserving surgery. However, the competing model rather contoured on the opposite breast, which was irrelevant to surgical method.
@@
Table 4: Ablation of input text information components for two different multimodal methods.
Numeric Category (Competing) LLMSeg (Ours)
Ablated Input Text Components Ablated Input Text Components
None Laterality N stagae T stage Surgery None Laterality N stagae T stage Surgery
Internal Test (N=307) 0.821 (0.805-0.836) 0.720 (0.695-0.744) 0.785 (0.768-0.800) 0.793 (0.777-0.808) 0.791 (0.771-0.809) 0.829 (0.809-0.845) 0.424 (0.383-0.467) 0.799 (0.781-0.815) 0.816 (0.795-0.832) 0.791 (0.771-0.809)
0.715 (0.697-0.733) 0.599 (0.573-0.624) 0.664 (0.646-0.680) 0.674 (0.657-0.691) 0.680 (0.659-0.699) 0.730 (0.709-0.748) 0.345 (0.309-0.382) 0.686 (0.667-0.702) 0.712 (0.690-0.730) 0.680 (0.659-0.699)
5.387 (4.699-6.129) 8.758 (7.874-9.715) 6.042 (5.346-6.802) 5.932 (5.198-6.707) 5.050 (4.531-5.583) 3.386 (2.890-3.949) 12.751 (11.894-13.530) 5.071 (4.570-5.646) 3.934 (3.415-4.522) 5.050 (4.531-5.583)
External Test #1 (N=206) 0.734 (0.710-0.755) 0.712 (0.690-0.734) 0.724 (0.695-0.749) 0.716 (0.694-0.738) 0.710 (0.677-0.740) 0.822 (0.805-0.836) 0.344 (0.305-0.385) 0.786 (0.773-0.798) 0.804 (0.787-0.820) 0.710 (0.677-0.740)
0.601 (0.577-0.623) 0.573 (0.551-0.596) 0.595 (0.567-0.620) 0.579 (0.557-0.603) 0.590 (0.559-0.619) 0.709 (0.689-0.727) 0.251 (0.219-0.283) 0.656 (0.641-0.671) 0.686 (0.666-0.705) 0.590 (0.559-0.619)
17.010 (15.829-18.165) 19.446 (18.102-20.765) 18.058 (16.906-19.214) 20.579 (19.284-21.768) 8.652 (7.724-9.616) 4.256 (3.471-5.176) 14.093 (13.197-14.991) 7.424 (6.450-8.481) 6.981 (5.946-8.065) 8.652 (7.724-9.616)
External Test #2 (N=205) 0.737 (0.717-0.755) 0.555 (0.513-0.595) 0.668 (0.638-0.694) 0.669 (0.638-0.698) 0.700 (0.681-0.718) 0.844 (0.826-0.857) 0.390 (0.340-0.434) 0.805 (0.790-0.817) 0.833 (0.816-0.846) 0.763 (0.740-0.784)
0.599 (0.577-0.618) 0.436 (0.401-0.469) 0.529 (0.502-0.555) 0.538 (0.507-0.567) 0.553 (0.534-0.573) 0.740 (0.722-0.756) 0.304 (0.263-0.341) 0.682 (0.667-0.696) 0.724 (0.706-0.739) 0.637 (0.613-0.660)
13.703 (12.706-14.724) 13.026 (12.144-13.812) 11.460 (10.562-12.343) 13.082 (11.945-14.242) 11.504 (10.538-12.383) 3.004 (2.555-3.533) 11.971 (11.161-12.814) 4.618 (4.244-5.029) 3.954 (3.475-4.475) 7.246 (6.571-7.951)
numeric category of each clinical data is ablated, e.g., “0301”, where the first 0 representing “N0” is replaced with a character ‘?’, resulting ‘?301”.
Exploring Other Cancer Types.
We further evaluated the proposed multimodal target volume contouring for prostate cancer patients, as shown in Table 5. Similar to the breast cancer study, we observed the superiority of our multimodal approach over the vision-only approach, with a notable performance gain of up to 0.05 in Dice metric through all the validation settings.
Similar to breast cancer, an expert evaluation was conducted for prostate cancer. A rubric-based analysis of expert evaluation in Table 6 clearly showed effectiveness of our method. Particularly, these benefits became unequivocally evident in the external validation setting, showing more than double the differences in total scores. Among those necessitating in-depth reference to clinical information for precise scoring — notably, the delineation of the primary site (assessing prostate volume coverage, including the seminal vesicle) and the volume definition (evaluating regional node irradiation appropriateness) — exhibited significantly larger differences compared to the vision-only model. Details on the rubrics used for prostate cancer can be found in the Supplementary Fig. S2 and Supplementary Table S2.
Table 5: Comparison of CTV delineation performance for prostate cancer patients.
@@
+0.047 +0.054 -0.240
Note. blue font denotes the performance difference between the vision-only AI and the multimodal AI
@@
Rubric Vision-only AI Multimodal AI (LLMSeg)

### v4

@@ Acknowledgement
Tuning Method Text prompts - Text prompts One text prompt LoRA [28] No tuning
@@
External Test #2 (N=204) Dice $\uparrow$ 0.844 (0.826-0.857) 0.737 (0.717-0.755) 0.528 (0.490-0.567) 0.832 (0.815-0.846) 0.796 (0.779-0.810) 0.829 (0.812-0.842)
@@
numeric categorization of text, e.g., “0301”, corresponds to “N0” for the first digit, ‘T3” for the second, “mastectomy” for third, and “right” for the last.
@@
Table 4: Ablation of input clinical data components for two different multimodal methods.
Ablated Input Text Components
Dataset No ablation Laterality N stage T stage Surgery
LLMSeg (Ours)
Internal Test (N=307) 0.829 (0.809-0.845) 0.424 (0.383-0.467) 0.799 (0.781-0.815) 0.816 (0.795-0.832) 0.791 (0.771-0.809)
0.730 (0.709-0.748) 0.345 (0.309-0.382) 0.686 (0.667-0.702) 0.712 (0.690-0.730) 0.680 (0.659-0.699)
3.386 (2.890-3.949) 12.751 (11.894-13.530) 5.071 (4.570-5.646) 3.934 (3.415-4.522) 5.050 (4.531-5.583)
External Test #1 (N=206) 0.822 (0.805-0.836) 0.344 (0.305-0.385) 0.786 (0.773-0.798) 0.804 (0.787-0.820) 0.710 (0.677-0.740)
0.709 (0.689-0.727) 0.251 (0.219-0.283) 0.656 (0.641-0.671) 0.686 (0.666-0.705) 0.590 (0.559-0.619)
4.256 (3.471-5.176) 14.093 (13.197-14.991) 7.424 (6.450-8.481) 6.981 (5.946-8.065) 8.652 (7.724-9.616)
External Test #2 (N=204) 0.844 (0.826-0.857) 0.390 (0.340-0.434) 0.805 (0.790-0.817) 0.833 (0.816-0.846) 0.763 (0.740-0.784)
0.740 (0.722-0.756) 0.304 (0.263-0.341) 0.682 (0.667-0.696) 0.724 (0.706-0.739) 0.637 (0.613-0.660)
3.004 (2.555-3.533) 11.971 (11.161-12.814) 4.618 (4.244-5.029) 3.954 (3.475-4.475) 7.246 (6.571-7.951)
Numeric Category (Competing method)
Internal Test (N=307) 0.821 (0.805-0.836) 0.720 (0.695-0.744) 0.785 (0.768-0.800) 0.793 (0.777-0.808) 0.791 (0.771-0.809)
0.715 (0.697-0.733) 0.599 (0.573-0.624) 0.664 (0.646-0.680) 0.674 (0.657-0.691) 0.680 (0.659-0.699)
5.387 (4.699-6.129) 8.758 (7.874-9.715) 6.042 (5.346-6.802) 5.932 (5.198-6.707) 5.050 (4.531-5.583)
External Test #1 (N=206) 0.734 (0.710-0.755) 0.712 (0.690-0.734) 0.724 (0.695-0.749) 0.716 (0.694-0.738) 0.710 (0.677-0.740)
0.601 (0.577-0.623) 0.573 (0.551-0.596) 0.595 (0.567-0.620) 0.579 (0.557-0.603) 0.590 (0.559-0.619)
17.010 (15.829-18.165) 19.446 (18.102-20.765) 18.058 (16.906-19.214) 20.579 (19.284-21.768) 8.652 (7.724-9.616)
External Test #2 (N=204) 0.737 (0.717-0.755) 0.555 (0.513-0.595) 0.668 (0.638-0.694) 0.669 (0.638-0.698) 0.700 (0.681-0.718)
0.599 (0.577-0.618) 0.436 (0.401-0.469) 0.529 (0.502-0.555) 0.538 (0.507-0.567) 0.553 (0.534-0.573)
13.703 (12.706-14.724) 13.026 (12.144-13.812) 11.460 (10.562-12.343) 13.082 (11.945-14.242) 11.504 (10.538-12.383)
for numeric categorization method, each digit is ablated, e.g., “0301”, where the first 0 representing “N0” is
replaced with a character ‘?’ to yield “?301”.
Table 5: Comparison of 3D CTV delineation performance for prostate cancer patients.
@@
Expert Rubrics